lilbunnybun:

If a x b = 2a - 3b + ab, then 3 x 5 + 5 x 3 is equal to? 20 22 24 26

xxemilianaxx:

what do you think

lilbunnybun:

26 but im not quite sure qwq

JamesTDG:

is that x a multiplication symbol or a variable?

lilbunnybun:

x

AZ:

\(\color{red}{a} \times \color{orange}{b} = 2\color{red}{a} - 3\color{orange}{b} + \color{red}{a}\color{orange}{b}\)

AZ:

So what is \(\color{red}{3} \times \color{orange}{5} = 2\color{red}{(3)} - 3\color{orange}{(5)} + \color{red}{(3)}\color{orange}{(5)}\) \(\color{red}{5} \times \color{orange}{3} = 2\color{red}{(5)} - 3\color{orange}{(3)} + \color{red}{(5)}\color{orange}{(3)}\)

AZ:

And then we add those two together

AZ:

So let's do the first one first \(\color{red}{3} \times \color{orange}{5} = 2\color{red}{(3)} - 3\color{orange}{(5)} + \color{red}{(3)}\color{orange}{(5)}\) What is 2 * 3 = ? What is 3 * 5 = ? What is 3 * 5 = ?